OpenStudy (anonymous):

Solve 5x + 14 - 2x = 9 - (4x + 2) a) x =3/7 b) x = - 1 c) x = 6 d) x = 8

OpenStudy (anonymous):

5x+14-2x=9-4x-2 5x-2x+14=-4x+9-2 3x+14=-4x+7 +4x-14+4x-14 7x=-7 x=-1

OpenStudy (theviper):

\[\LARGE{\color{red}{3x+14=7-4x}}\]\[\LARGE{\color{green}{\implies4x+3x=7-14 }}\]\[\LARGE{\color{green}{\implies 7x=-7}}\]\[\LARGE{\color{green}{\implies x=-1}}\]

OpenStudy (anonymous):

Thank you!

OpenStudy (theviper):

\[\Huge{\mathbb{\text{Got it??}}}\]

OpenStudy (anonymous):

Yes Randy Orton lol

OpenStudy (theviper):

I think u have got it that's why u said \[\Huge{\color{gold}{\star \star \star}}\]Thank u!